OpenStudy (jkbo):

The perimeter of a rectangle with base 10 in. and height 13 in. is _____. Be sure to include units. (1 point

OpenStudy (anonymous):

2(10+13) 2(23) 46in.

OpenStudy (anonymous):

P = 2(base + height) P = 2(10 +13) P = 2(23) P = 46 So the perimeter is 46 inches. (same thing @monokerous got)

OpenStudy (jkbo):

thanks

OpenStudy (anonymous):

np (: